OpenStudy (anonymous):

how do i draw a vector function of <3sint, cost> ?

OpenStudy (anonymous):

are you using MATLAB?

OpenStudy (anonymous):

what's MATLAB?

OpenStudy (anonymous):

if your in college you must be knowing that, its a software for solving everything :) known as MATRIX lab or matlab and your vector function could have been drawn on that easily

OpenStudy (anonymous):

Oh. well this is my first math class since high school, and im having trouble remembering everything and my book isnt doing me much justice..so how do i get there?

OpenStudy (anonymous):

i guess phasor diagram can be plotted of the function that you have given but vector function is kinda complex :/ to draw , you gotta use software

OpenStudy (anonymous):

but i have to draw it on paper...

OpenStudy (anonymous):

are you a math major?

OpenStudy (anonymous):

supposedly...now im not so sure...i cant even comprehend calc 3

OpenStudy (anonymous):

don't worry you can go through that i've forgotten this, lemme see how much i can do here :/

OpenStudy (anonymous):

okay thanks! like i dont even know where to begin..

OpenStudy (anonymous):

got it, just use different values of t and put it in your function, you can draw vector function easily, lets say at t=30 degrees ,find value of 3 sint which is = 3 * 1/2 and value of cos t = cos 30 = sqrt 3/ 2 so you got a pair that is \[(\frac{ 3 }{ 2 }, \frac{ \sqrt 3 }{ 2 } )\] similarly find value of your function at lets say t=45 your gona get 3 sin t = 3 sin 45 = 3/ sqrt 2 and cos 45 = 1/ sqrt 2 \[(\frac{ 3 }{ \sqrt2 }, \frac{ 1 }{ \sqrt2 })\] now just plot it on graph paper ;)

OpenStudy (anonymous):

did you get this ? @destinie

OpenStudy (anonymous):

yes im trying to decipher and make sense of it but thank you sooooo much!

OpenStudy (anonymous):

yeah i think i understand how to plot them, but since they are a vector function, isnt it supposed to be a vector? or do i plot them and then thats the vector? and i attend IUS lol nice to meet you!

OpenStudy (anonymous):

 the line that i have shown by arrow is the only line just find out values at t= pi/4, pi/2 and pi finally you are done ;)